# skipabove=1em,skipbelow=1em

### v3

@@ 7.2.1. Modality Misalignment between Sensing and Intervention Phases
Sensing studies have made significant strides in utilizing a wide array of modalities, including wearable devices, ambient sensors, and smartphone-integrated tools, to detect ingestive behaviors with increasing granularity and accuracy [e.g., (Bedri et al., 2022; Zhang et al., 2020)]. These tools offer real-time monitoring of behaviors like food intake, bite frequency, and meal duration. In contrast, interventions rely on a narrower set of modalities, mostly using visual cues like text prompts, graphical feedback, or auditory signals delivered via mobile devices. Although some studies [e.g., (Kim et al., 2016b)] have used wearable devices for both sensing and delivering interventions, such integrated applications remain relatively scarce, as shown in Chapter 5.3. This disconnect between detection and intervention modalities creates a lack of continuity in the intervention process, limiting the potential for closed-loop systems where detection and feedback are fully integrated. For instance, while detection technologies capture behavioral data, the feedback often remains confined to basic prompts, which may not effectively engage users or encourage sustained behavior change.
@@ 7.2.2. Disconnect Between Detection-Centric Studies and Actual Intervention Implementation
A second significant gap exists in the large body of research only focus on improving detection accuracy ,but without progressing to actionable interventions. Many sensor-based studies mention behavior change as a long-term goal but fall short of translating their findings into real-world applications. This disconnection is largely due to the differing complexities of sensing and intervention research. Detection-centric studies typically focus on technical performance metrics, such as the accuracy of behavior detection and system usability, which are easier to evaluate. On contrary, intervention studies must deal with a broader range of challenges, including incorporating behavior change theories, designing meaningful user interactions, and establishing long-term engagement. The higher threshold for evaluating interventions has contributed to a bottleneck, where many studies stop after validating detection methods, while fewer studies advance to actual intervention implementation. As a result, there is an imbalance in the literature, with a higher volume of studies dedicated to detection accuracy, and fewer addressing the complexity of behavior modification and sustained interventions.
@@
Although both Human-Computer Interaction (HCI) and medical research aim to improve ingestive behaviors, such as dietary patterns and meal composition, they differ in their assessment frameworks and metrics. HCI typically focuses on real-time, sensor-based interventions and user experience, whereas medical research emphasizes clinical assessments and physiological measures. These differences pose challenges in aligning the methodologies and measuring the effectiveness of intervention. Additionally, while both fields acknowledge the importance of subjective experiences and physiological states, their integration into behavior metrics differs. HCI research tends to incorporate subjective feedback via interactive devices, while medical research relies more heavily on objective physiological data. This gap highlights the limitations of current HCI technologies in capturing the complex physiological aspects of ingestive behavior.
@@ 7.3.1. Divergent Assessment Frameworks
From a medical perspective, healthy ingestive behaviors are often linked to specific health conditions, such as cardiovascular disease, type 2 diabetes, obesity, and cancer. Medical practitioners use comprehensive nutritional assessments that combine food history, biochemical data, medical tests, and clinical procedures (Kesari and Noel, 2022). Tools like the Food Frequency Questionnaire (FFQ) (Cade et al., 2004), Three-Factor Eating Questionnaire (TFEQ) (Karmali et al., 2013), and Adult Eating Behavior Questionnaire (AEBQ) (Hunot et al., 2016) offer a validated, multi-domain approach for assessing ingestive behaviors and their relationship to health outcomes.
In contrast, HCI research often relies on isolated validation methods, lacking the cross-validation mechanisms inherent in medical assessments. While HCI interventions focus on sensor-based technologies and user behavior tracking, they do not integrate the multi-domain assessments seen in medical protocols. This difference limits HCI-based interventions in capturing the complex interactions between physiological states and behavioral patterns. HCI tools primarily focus on real-time data acquisition (e.g., food intake, eating speed), but they often overlook broader health contexts, such as digestive symptoms or stress levels. This results in an incomplete understanding of the long-term impact of HCI interventions on health outcomes.
@@ 7.3.2. Limited Perception of Subjective Experience and Physiological States in Target Behavior Metrics
Medical assessments integrate subjective experiences (e.g., emotional eating, satiety cues) with physiological states, offering a comprehensive evaluation of ingestive behaviors essential for disease diagnosis (Kesari and Noel, 2022). These tools benefit from (1) long-term validation, (2) structured clinical environments that enhance patient compliance, and (3) professional interpretation by healthcare specialists. These assessments are particularly effective in documenting longitudinal changes in variables such as pain, function, and psychosocial distress.
In contrast, early HCI research relied on questionnaires that (1) had limited long-term validation, (2) faced low accuracy and compliance due to experimental constraints, and (3) lacked professional interpretation. While HCI interventions track behaviors like chewing or swallowing, they fail to capture subjective experiences (e.g., satiety or emotional states), essential for understanding ingestive behaviors comprehensively. This gap highlights a key limitation: sensors can tell us how we eat, but they do not address why we eat or how we feel during eating, which are crucial aspects for understanding ingestive behaviors in the context of health outcomes.
Thus, while HCI tools focus on quantifiable metrics (e.g., food intake or eating speed), they overlook the subjective aspects emphasized by medical assessments. To address this gap, HCI interventions should integrate both subjective and physiological data to offer a more holistic understanding of ingestive behaviors, particularly in real-world contexts.
@@ 7.4.1. Optimizing Experimental Design and Intervention Implementation
To bridge the gap between laboratory-based research and real-world interventions, clear and standardized measurement metrics should be established, drawing from validated constructs in the existing literature. Combining qualitative and quantitative methods provides a comprehensive view of intervention outcomes, helping reduce bias and ensuring a more accurate evaluation. Additionally, integrating self-reported data with objective physiological measures offers a balanced approach to long-term tracking, avoiding participant fatigue. For example, targeted micro-assessments, triggered by unusual behavior patterns, can provide timely and context-specific insights without overwhelming participants.
When designing experimental protocols, it is essential to consider real-world conditions, user experience, and long-term feasibility. Instead of excluding common unhealthy behaviors, such as continuous eating without proper pauses, interventions should directly address them as part of the behavior change process. This ensures that interventions are more relevant to natural eating contexts, promoting sustainable behavior change.
For long-term engagement, the technical design should prioritize usability and integration with daily life. Embedding sensing capabilities into familiar, unobtrusive devices—such as smartwatches with inertial sensors or bone conduction earphones—can enhance natural interaction and reduce disruption. This approach will ensure continuous monitoring while preserving comfort and user adherence.
Furthermore, ethical safeguards must be incorporated into the intervention design. Participants should be clearly informed that they can withdraw at any time if they feel discomfort or emotional distress, ensuring their psychological well-being is not compromised by the intervention.
Finally, given the inconsistency in measurement tools across studies, future interventions should consider context-sensitive evaluation methods. These should account for cultural differences and ensure interventions align with both natural ingestive behaviors and social norms, improving applicability across diverse populations.
@@ 7.4.2. Enhancing Integration and Adaptability in Sensing-Driven Interventions for Ingestion Health
To effectively bridge the gap between sensing and intervention in ingestion health, future interventions should focus on the seamless integration of real-time data from sensing technologies into personalized, actionable, and adaptive interventions. To enhance the continuity of the closed-loop system, interventions should increasingly leverage the same or compatible devices for both sensing and feedback delivery. Such consistency in device use not only improves the intuitiveness and coherence of the intervention process but also fosters user familiarity and acceptance over time. More personalized and actionable intervention requires moving beyond basic detection capabilities, not only by gathering more granular physiological and behavioral data but also by ensuring that feedback mechanisms are tailored to the specific needs, preferences, and circumstances of each user. For instance, interventions must consider factors such as individual eating patterns, meal timing, emotional states, and physiological markers like glucose levels, enabling a level of personalization that ensures more effective engagement and results (Liang, 2024; Reynolds et al., 2019).
Moreover, interventions should be designed with adaptability in mind. Success in behavior change depends not only on tracking short-term ingestive behaviors but also on continually adapting the intervention based on long-term data. This means creating systems that can evolve their feedback strategies in response to behavioral trends and progress over time. Such adaptive interventions maintain their relevance and motivational power, preventing user fatigue and promoting long-term engagement. A crucial aspect of this adaptability involves understanding the dynamic nature of ingestive behaviors. For example, certain behaviors may be triggered by emotional states or specific environmental factors, which require intervention systems that can quickly detect and respond to these variables. This requires the continuous collection and analysis of both physiological and behavioral data, ensuring that interventions are timely and appropriate.
Finally, the long-term success of such systems hinges on the ability to sustain behavior change. The intervention should encourage not just short-term compliance but foster lasting habits by gradually increasing the complexity or frequency of the feedback, thereby encouraging the user to internalize healthier behaviors and develop autonomous control over their eating habits. This continuous, evolving support through adaptive interventions will ensure that users are not only supported in their current behavior change but are also equipped with the tools to sustain these changes in the long run.
@@ 7.4.3. Aligning with Medical Goals and Practices
To bridge the gap between HCI-based interventions and medical objectives, personalized interventions should account for the diversity in ingestive behaviors, physiological baselines, and cultural backgrounds among participants (Mirmiran et al., 2021). Incorporating both subjective and physiological data, especially baseline parameters like blood glucose and heart rate variability, will improve the precision and effectiveness of interventions (Liang, 2024; Harous et al., 2017).
Future studies should adopt dynamic intervention models and extend trial durations to better capture long-term effects. Exploring dose-response relationships, including varying intervention frequency, duration, and intensity, will help identify the optimal parameters for sustained health benefits (Pankajavalli et al., 2017; Kim et al., 2016b).
Emerging behavioral and physiological targets such as postprandial heart rate variability, blood glucose response curves, and digestive symptoms (e.g., bloating, nausea) offer valuable insights into metabolic health. Continuous glucose monitoring (CGM) has shown promise in real-time glucose tracking, and integrating it with wearable technologies, including smartwatches and fitness trackers, can offer multi-dimensional feedback on ingestive behavior and metabolic response (Liang, 2024; Bedri et al., 2022; Zhao et al., 2021).
Multimodal sensing approaches, integrating wearable and minimally invasive technologies, could enable more comprehensive tracking of physiological and behavioral signals, enhancing intervention outcomes (Zhang and Parker, 2020). Future sensor fusion techniques, combining CGM with emotion detection (EDA) and activity tracking, will offer a complete view of ingestive behavior and its impact on health (Carroll et al., 2013).
As sensor technology progresses, real-time systems could detect subtle digestive processes, like esophageal peristalsis during swallowing, to assess intervention effects on digestive function (Tirasirichai et al., 2018; Biyani et al., 2019). These advancements will improve our understanding of nutrition-physiology relationships and provide actionable insights for chronic disease prevention (Reynolds et al., 2019; Bi et al., 2016).
@@ 8. Conclusion and Limitations
In this review, we systematically analyzed 136 studies focused on behavioral sensing and intervention paradigms in ingestion health, with an emphasis on sensor-based systems and interactive interventions. These studies were categorized into various types based on the modality of feedback and the behavioral targets of the interventions. A key contribution of this review is the development of a behavioral closed-loop framework, which highlights the importance of seamlessly integrating sensing, reasoning, and feedback mechanisms for effective behavior change. This framework offers a detailed structure for the design of adaptive, context-aware interventions, targeting both physiological and behavioral aspects of ingestive health.
However, despite the comprehensive nature of this review, there are several limitations that should be acknowledged. Our keyword search strategy, centered around the term “ingestion health,” may have excluded relevant studies that do not explicitly focus on ingestion but contribute to broader research in behavior modification or health interventions. Moreover, the majority of the studies in HCI domain reviewed were short-term, laboratory-based experiments, which limits the generalizability of our findings to real-world settings. The dynamic nature of real-life eating behaviors requires more longitudinal studies to better understand the long-term effectiveness and challenges of these interventions.

### v4

@@ 7.2.1. Modality Misalignment between Sensing and Intervention Phases
Sensing studies have employed diverse modalities, including wearable devices, ambient sensors, and smartphone-integrated tools, to detect ingestive behaviors with increasing granularity and accuracy [e.g., (Bedri et al., 2022; Zhang et al., 2020)]. These systems offer real-time monitoring of behaviors like food intake, bite frequency, and meal duration. In contrast, most intervention applications rely on a narrow set of output channels, typically visual cues like text prompts, graphical feedback, or auditory signals via mobile devices. Although some efforts [e.g., (Kim et al., 2016b)] have explored using the same device for both sensing and intervention, such integrated applications remain uncommon, as shown in Chapter 5.3. This misalignment between sensing and intervention modalities disrupts continuity in the closed-loop process — while sensors capture detailed, real-time behavioral data, interventions are often reduced to basic prompts that struggle to engage users or support long-term behavior change. This gap calls for tighter integration between sensing and intervention modalities through compatible devices, enabling a unified user experience.
@@ 7.2.2. Disconnect Between Detection-Centric Studies and Actual Intervention Implementation
Many sensor-based studies of ingestive behavior emphasize detecting and monitoring intake-related events but stop short of delivering actionable interventions. While behavior change is often cited as a long-term goal and occasionally guided by behavioral frameworks (Hekler et al., 2013), most studies ultimately focus on improving detection accuracy, as noted in Section 5. This imbalance stems from differing evaluation paradigms: detection studies prioritize quantifiable metrics (e.g., recognition accuracy), which are easier to validate in controlled settings, whereas interventions are often treated as downstream demonstrations—supporting sensing pipelines rather than contributing novel design knowledge (Rapp and Boldi, 2023), as discussed in Chapter 6. Designing interventions, however, involves more complex challenges, including applying behavior change theories, supporting meaningful user interaction, and sustaining long-term engagement—factors difficult to capture through standard evaluation methods. These demands contribute to a bottleneck where sensing advances outpace the development of effective intervention strategies.
Beyond methodological barriers, a deeper issue lies in the limited integration of behavioral theory. Although behavior theories are often cited, they rarely inform system design in a structured way (Hekler et al., 2013). For instance, the Fogg Behavior Model (Fogg, 2009) defines behavior as a convergence of motivation, ability, and trigger. Though some interventions loosely map sensing accuracy to ”ability”, context to ”motivation”, and reminders to ”triggers”, often retrofitted to justify design choices. In our review, few systems explicitly used such frameworks to shape intervention logic or personalize support. This shallow application reflects a broader reliance on intuition over theory, with key constructs, like habit, emotion, and social support, frequently overlooked. As a result, interventions may be theoretically underpowered and less effective in driving sustained behavior change.
@@
Although both Human-Computer Interaction (HCI) and medical research aim to promote healthier ingestive behaviors, they differ significantly in how they define, measure, and interpret such behaviors. HCI approaches tend to focus on context-aware computing and interactive methods (Nahum-Shani et al., 2016; Hekler et al., 2016), while medical research emphasizes clinically validated assessments and physiological indicators. These differences pose challenges in aligning methodologies, interpreting efficacy, and integrating subjective experiences and biological states. In the following sections, we explore two key dimensions of this divergence — the assessment frameworks and treatment of subjective and physiological aspects of ingestive behavior.
@@ 7.3.1. Divergent Assessment Frameworks
From a medical perspective, healthy ingestive behaviors are often linked to specific health conditions, such as cardiovascular disease, type 2 diabetes, obesity, and cancer. Medical practitioners use comprehensive nutritional assessments that combine food history, biochemical data, medical tests, and clinical procedures (Kesari and Noel, 2022). For instance, validated instruments such as the Food Frequency Questionnaire (FFQ) (Cade et al., 2004), Three-Factor Eating Questionnaire (TFEQ) (Karmali et al., 2013), and Adult Eating Behavior Questionnaire (AEBQ) (Hunot et al., 2016) can capture multiple aspects, like appetite traits, restraint, and emotional eating, and help relate ingestion patterns to risks.
In contrast, HCI research often relies on isolated validation methods, lacking the multi-domain cross-validation seen in clinical practice. For example, HCI studies rarely combine physiological indicators (e.g., glucose, hormones) with psychological measures to evaluate intervention effects. This limits the ability to capture interactions between physiological states and behavioral patterns. HCI tools focus on real-time data (e.g., food intake, eating speed), but often ignore broader health factors like stress or digestion. As a result, their long-term health impact remains unclear. Medical approaches, in contrast, emphasize multi-dimensional and longitudinal assessments, offering a deeper view of intervention effectiveness.
@@ 7.3.2. Limited Perception of Subjective Experience and Physiological States in Target Behavior Metrics
Medical assessments emphasize the integration of subjective experiences (e.g., emotional eating, satiety, post-meal discomfort) with physiological states to form a comprehensive evaluation of ingestive behavior, particularly for diagnosing and managing conditions like functional dyspepsia or eating disorders (Kesari and Noel, 2022). These assessments benefit from validated instruments with long-term reliability, structured clinical environments that support patient compliance, and professional interpretation by healthcare practitioners. For example, longitudinal evaluation of chronic conditions often relies on self-reported changes in symptoms (e.g., pain, bloating, appetite) combined with physiological data to guide treatment.
By contrast, most HCI studies rely on short-term self-report tools with limited validation and low compliance in non-clinical settings. While sensor-based systems excel at quantifying behaviors (e.g., eating rate, chewing duration), they often fail to capture the subjective dimensions essential to understanding why users eat the way they do. Early HCI studies seldom included validated measures of satiety or emotional states, and even with the rise of ubiquitous computing, the capacity to sense nuanced experiential factors remains limited. This overemphasis on observable behavior, without accounting for users’ internal states, restricts the interpretability and impact of HCI interventions.
@@ 7.4. Design Recommendations
To address the core challenges identified above, we propose design recommendations and future directions to support more effective and ecologically embedded closed-loop interventions for ingestive behavior. Advancing this domain requires coordinated efforts across researchers, designers, and practitioners, spanning theory, technical implementation, and real-world integration. The following subsections align our suggestions with the key challenges, offering concrete guidelines for each domain.
@@ 7.4.1. Optimizing Experimental Design and Intervention Implementation
To move beyond proof-of-concept prototypes toward impactful interventions, researchers should adopt standardized, theory-driven, and context-sensitive evaluation frameworks. Researchers should combine validated behavioral constructs with multimodal sensing data (e.g., self-report diaries and sensors) to contextualize outcomes and reduce measurement bias. Methodologically, designs such as factorial experiments, SMART, and micro-randomized trials are well-suited to evaluating multi-component systems (Nahum-Shani et al., 2022). Such designs enable precise testing of when and how interventions are most effective, ultimately informing evidence-based guidelines. To reduce burden while preserving ecological validity, researchers may adopt micro-assessments triggered by behavioral anomalies to gather timely and lightweight feedback. Grounding evaluations in behavioral theories helps bridge the disconnect between technological innovation and behavior change science.
Intervention tools should be intuitive, low-burden, and compatible with daily routines, such as embedding this loop into smart utensils or wearables. Co-locating sensing and intervention delivery within a single ecosystem improves usability and streamlines interaction. Designers should provide timely, actionable interventions rather than passive data to support real-time engagement. For example, FitNibble (Bedri et al., 2022) and Slowee (Kim et al., 2016b) demonstrate how subtle prompts can improve adherence and engagement in natural settings. Rather than filtering out common unhealthy behaviors in study design, interventions should address them explicitly as change targets to better reflect real-world eating challenges.
Privacy and trust should be core design priorities in always-on, closed-loop systems. Designers should implement privacy-preserving strategies, including context-aware data collection (Nepal et al., 2024; Saha et al., 2021) and layered encryption (Kumar et al., 2020). Practitioners should ensure transparent communication about data use and obtain sustained informed consent (Haghighi et al., 2023; Adler et al., 2022; Islam and Bae, 2024). Systems should enable user autonomy, allowing for flexible participation and empathetic adjustments when engagement decreases or discomfort arises.
To enhance generalizability, practitioners should iteratively adapt interventions to diverse physiological and cultural contexts. Integrating behavioral coaching with tech-driven systems ensures more personalized and effective support. By combining scientific rigor, user-centered design, and ethical awareness, interventions can better support sustained ingestive behavior change in the real world.
@@ 7.4.2. Enhancing Integration and Adaptability in Sensing-Driven Interventions for Ingestion Health
To close the gap between sensing and intervention in ingestion health, future systems must better integrate behavioral theory into personalized, adaptive strategies. Many recent interventions emphasize surface-level mechanisms, like rational reflection or automated nudges, without deeply engaging with why those mechanisms work across individuals. This limits their long-term impact, as users may revert to old habits once external support ends.
Researchers should embed theory-driven constructs into system logic. Behavior change theories such as the Fogg Behavior Model (Fogg, 2009) or Social Cognitive Theory offer constructs like motivation, ability, and self-efficacy that should guide not just evaluation but delivery decisions — when to intervene, what content to present, and how to adapt dynamically. Rather than cherry-picking isolated elements, researchers should operationalize full frameworks to understand both success and failure. They should also personalize adaptively. Building on the JITAI paradigm (Nahum-Shani et al., 2016), systems can adjust message type and timing based on real-time behavioral or physiological trends. For instance, if late-night snacking consistently correlates with stress spikes, systems may prioritize emotion regulation prompts over dietary tracking.
Designers should tightly couple sensing and intervention in a closed loop. Physiological signals like chewing rate or glucose dips should directly trigger timely, context-aware interventions. Integrated systems, such as SenseSupport (Liang, 2024), reduce user friction and enhance moment-to-moment relevance. They should also scaffold long-term autonomy. Systems should evolve with the user—starting with frequent, specific prompts, then gradually tapering or increasing complexity to foster independence.
Practitioners should ground behavior change in lived experience. Traditional models often abstract behavior from context, but effective change requires alignment with users’ identities, values, and goals (Lee et al., 2020). To foster sustained motivation, interventions should facilitate meaning-making by helping users relate new eating habits to their personal values, such as family responsibility or personal development (Rapp and Boldi, 2023).
@@ 7.4.3. Aligning with Medical Goals and Practices
For sensing-driven interventions to achieve lasting health benefits and clinical relevance, they must align with medically meaningful outcomes and practices. Researchers should work with health professionals to define targets beyond behavioral proxies—such as improvements in glycemic control, cholesterol, gastrointestinal symptoms, or long-term weight trends. Incorporating baseline physiological parameters (e.g., glucose levels, heart rate variability) alongside behavioral data can enhance personalization and clinical interpretability (Liang, 2024; Harous et al., 2017; Mirmiran et al., 2021).
Future studies should adopt longer-term and dynamic intervention models to assess sustained physiological effects. Exploring dose-response relationships and optimizing frequency, intensity, and timing can help identify scalable strategies (Pankajavalli et al., 2017; Kim et al., 2016b). Drawing on techniques from clinical behavior change approaches—such as motivational interviewing or cognitive-behavioral therapy—can further ground interventions in established medical frameworks.
Multimodal sensing and sensor fusion offer promising directions. Combining continuous glucose monitoring (CGM), emotion sensing (e.g., EDA), activity tracking, and digestive symptom monitoring enables more comprehensive understanding of ingestive behavior and metabolic response (Bedri et al., 2022; Zhang and Parker, 2020; Zhao et al., 2021). These data streams can support real-time personalization, such as prompting slower eating when postprandial glucose rises or initiating mindfulness exercises during periods of physiological stress.
Finally, aligning with medical goals invites broader thinking about precision health. Though this work centers on eating behavior, the underlying strategies can extend to domains like sleep, medication adherence, and stress management. Framing these systems within the larger precision health movement will improve scalability, adoption, and long-term public health impact.
7.4.4. Extending the Paradigm Beyond Ingestion Health
The proposed behavioral closed-loop paradigm, while developed in the context of ingestion health, offers a generalizable framework for a wide range of behavior change interventions. As demonstrated by prior reviews (Lara and Labrador, 2012; Thomas Craig et al., 2021; Mair et al., 2023), wearable and mobile sensing technologies have been successfully employed in domains such as physical activity promotion, dietary behavior optimization, chronic disease management, substance use reduction, and mental health support. These systems often combine continuous sensing (e.g., step counts, glucose levels, stress markers), reasoning components (e.g., activity inference, risk prediction), and real-time interventions (e.g., feedback messages, prompts), aligning closely with the structure of our proposed closed-loop model.
Continuous, context-aware sensing should serve as the foundation for detecting actionable behavior patterns across domains. For instance, physical activity interventions detect sedentary behavior and deliver timely prompts to encourage movement (Boerema et al., 2019; Van Dantzig et al., 2018); chronic disease management systems combine physiological sensing with persuasive feedback to improve self-management (Chandler et al., 2019; Coorey et al., 2019). Real-time, personalized intervention strategies should adapt to users’ physiological, psychological, and contextual states. These examples highlight the potential to extend our framework beyond eating-related behaviors while retaining its core responsiveness and personalization.
The closed-loop structure provides a modular blueprint that can unify design across domains, enabling cross-disciplinary translation and theory-driven development. By abstracting core design elements — continuous sensing, adaptive reasoning, and context-aware intervention — the paradigm facilitates broader scalability and long-term health impact in ubiquitous computing.
@@ 8. Conclusion and Limitations
In this review, we systematically analyzed 136 studies focused on behavioral sensing and intervention paradigms in ingestion health, with an emphasis on sensor-based systems and interactive interventions. These studies were categorized into various types based on the modality and the behavioral targets of the interventions. A key contribution of this review is the development of a behavioral closed-loop framework, which highlights the importance of seamlessly integrating sensing, reasoning, and intervention mechanisms for effective behavior change. This framework offers a detailed structure for the design of adaptive, context-aware interventions, targeting both physiological and behavioral aspects of ingestion health. It is theoretically grounded in context-aware computing and HCI behavior change frameworks, offering a modular foundation for system design across domains.
However, despite the comprehensive nature of this review, there are several limitations that should be acknowledged. Our keyword search strategy, centered around the term ”ingestion health”, may have excluded relevant studies that do not explicitly focus on ingestion but contribute to broader research in behavior modification or health interventions. Moreover, the majority of the studies in HCI domain reviewed were short-term, laboratory-based experiments, which limits the generalizability of our findings to real-world settings. The dynamic nature of real-life eating behaviors requires more longitudinal studies to better understand the long-term effectiveness and challenges of these interventions.
Finally, the modular structure of the proposed paradigm — grounded in continuous sensing, adaptive reasoning, and real-time intervention — holds promise for application beyond ingestion health, such as in sleep, stress, or physical activity interventions. We hope this review helps lay the foundation for more generalizable, stakeholder-sensitive, and theoretically informed systems in ubiquitous health computing.